OpenStudy (anonymous):

Hi, If cos(a+b)=cosacosb-sinasinb and if cos(a- b)=cosacosb+sinasinb then cosacosb=?

OpenStudy (anonymous):

I think the answer is like 1/2[something] but I dont know where to start!

OpenStudy (anonymous):

Has to do with trig formulas but I need some help.

jhonyy9 (jhonyy9):

so if cos(a+b)=cosacosb -sinasinb and cos(a-b)=cosacosb +sinasinb so than if you assum these will get ------------------ cos(a+b +cos(a-b)=2cosacosb and now just divide both sides by 2 and will get cos(a+b)+cos(a-b) ----------------- = cosacosb 2 hope so much that is understandably sure rigth good luck bye